# Generative Neural Video Compression via Video Diffusion Prior

### v1

@@ Abstract
Figure 1: Qualitative comparison on ultra-low bitrate video compression. Traditional and learned codecs produce blurry frames. Generative approaches such as GLC-Video [qi2025generative] yield sharper textures but introduce structural hallucinations and unstable details, causing pronounced temporal flickering (see Fig. 2). Leveraging a video-native diffusion prior, GNVC-VD produces coherent fine textures with strong temporal stability. Zoom in for best view.
@@ 1 Introduction
Neural video compression (NVC) [DVC, FVC, DCVC, DCVC-TCM, DCVC-HEM, DCVC-DC, DCVC-FM, DCVC-RT] has advanced rapidly in recent years, with learned codecs now surpassing traditional hybrid standards such as HEVC [sullivan2012overview] and VVC [bross2021overview] in rate–distortion optimization. However, when bitrate drops to the ultra-low regime, distortion-driven objectives (e.g., MSE) tend to oversmooth textures and erase fine structures, causing a sharp decline in perceptual realism. Improving perceptual quality under such extreme compression remains an open and fundamental challenge for NVC.
@@
In the image domain, this perceptual collapse has been largely alleviated. Recent generative image codecs [mao2024extreme, xue2024unifying, jia2024generative, careil2023towards, RDEIC, muckley2023improving, zhang2025stablecodec, ke2025ultra] leverage large pre-trained GANs [gan, vqgan] or diffusion models [stablediffusion] to recover high-frequency textures, producing visually convincing reconstructions even at extremely low bitrates. This naturally raises the question: can the same strategy be extended to video compression?
@@
Unfortunately, videos impose a much stricter requirement—temporal coherence. Although recent perceptual video codecs [ma2025diffusion, qi2025generative] integrate image generative priors [vqgan, stablediffusion], such priors remain inherently static and lack any modeling of temporal dynamics. As a result, codecs built upon them remain fundamentally frame-level: even with adjacent-frame conditioning, the generative prior cannot capture long-range temporal structure. Consequently, the restored appearance drifts over time, leading to the well-known perceptual flickering that becomes especially severe at ultra-low bitrates, as illustrated in Fig. 2.
@@
Recently, video diffusion models (VDMs)—especially those based on diffusion transformers (DiTs) [liu2024sora, yang2024cogvideox, kong2024hunyuanvideo, wan2025wan]—offer a natural path forward. Trained on large-scale video data, they learn spatio-temporal latent representations that capture appearance, motion, and long-range dependencies within a unified structure, enabling the synthesis of sequences with coherent texture and motion. These properties make VDMs an ideal generative prior for video compression, motivating us to rethink decoding not as independent frame reconstruction but as a sequence-level conditional denoising process guided by a video-native model.
@@
Building on this insight, we introduce GNVC-VD, the first generative NVC framework that fully leverages a pre-trained video diffusion transformer (VideoDiT). Unlike prior perceptual codecs [ma2025diffusion] constrained by image generative priors and thus limited to frame-wise enhancement, GNVC-VD redesigns the entire coding pipeline around sequence-level compression and generative refinement, enabling the diffusion prior to guide reconstruction beyond frame-wise prediction. At its core, GNVC-VD integrates two tightly coupled components: (1) a conditional contextual transform codec that compresses the spatio-temporal latent representations while preserving long-range temporal structure, and (2) a flow-matching–based latent refinement module that performs sequence-level generative denoising across both intra- and inter-frame latents, driven by the video DiT.
@@
Figure 2: (a) Spatial and ttxx comparisons. Traditional and learned codecs lose fine textures, while GLC-Video [qi2025generative] exhibits sharp but unstable structures that cause temporal flickering. GNVC-VD preserves clean textures and stable motion. (b) Frame-wise warp error $E_{\text{warp}}$ further confirms GNVC-VD’s temporal stability, in contrast to the large fluctuations of GLC-Video.
@@ 2 Related Works
Neural Video Compression [DVC, FVC, rippel2019learned, habibian2019video, DCVC, DCVC-TCM, DCVC-HEM, DCVC-DC, DCVC-FM, DCVC-RT, ma2025diffusion, qi2025generative] has made substantial progress in recent years, achieving strong RD performance across a wide range of bitrates. By learning compact latent representations and temporal dependencies in an end-to-end manner, NVC methods outperform traditional hybrid codecs in both PSNR and MS-SSIM metrics. Prior works can be broadly categorized into residual-based [DVC, FVC], 3D autoencoder-based [rippel2019learned, habibian2019video], and conditional coding-based architectures [DCVC, DCVC-TCM, DCVC-HEM, DCVC-DC, DCVC-FM, DCVC-RT]. Among these, conditional coding approaches—such as the DCVC family [DCVC, DCVC-TCM, DCVC-HEM, DCVC-DC, DCVC-FM, DCVC-RT]—have set new state-of-the-art results by using decoded features as context to guide motion estimation, latent prediction, and entropy modeling. However, as these methods are primarily optimized for distortion-oriented objectives (e.g., MSE), their reconstructions tend to be overly smooth and lack fine textures at extremely low bitrates. This reveals the inherent limitation of current NVC frameworks and calls for perceptually optimized compression to better preserve visual realism under extreme bitrate constraints.
@@
Perceptual Compression with Generative Prior. To enhance perceptual quality at ultra-low bitrates, recent studies introduce generative priors—termed generative compression [agustsson2019generative, mentzer2020high, muckley2023improving, mao2024extreme, xue2024unifying, jia2024generative, lei2023text+, careil2023towards, RDEIC, li2024towards, zhang2025stablecodec, ke2025ultra, qi2025generative, ma2025diffusion]—which leverage learned generative models to guide reconstruction and recover realistic textures beyond pixel fidelity. In the image domain, early works incorporate adversarial losses [agustsson2019generative, mentzer2020high, muckley2023improving] or VQ-based tokenization [mao2024extreme, xue2024unifying, jia2024generative] to achieve perceptually convincing reconstructions under extreme rate constraints. Diffusion-based frameworks [lei2023text+, careil2023towards, RDEIC, li2024towards, zhang2025stablecodec, ke2025ultra] further improve perceptual quality by reformulating decoding as conditional denoising guided by compact latents and leveraging large-scale text-to-image diffusion models prior [stablediffusion].
@@
Extending this paradigm to videos, recent methods such as GLC-Video [qi2025generative] and DiffVC [ma2025diffusion] adapt pre-trained image generative models by either encoding frame-wise generative embeddings or applying diffusion-based frame enhancement. However, these approaches still depend on image-domain priors without explicit temporal modeling, leading to flickering and motion inconsistency at ultra-low bitrates, as illustrated in Fig. 2(b). In contrast, our work is the first to introduce a video generative prior into NVC, jointly encoding spatio-temporal latents and performing unified latent refinement to achieve temporally coherent and perceptually realistic reconstructions under extreme compression.
@@
Video Diffusion Models have emerged as powerful generative frameworks capable of synthesizing high-quality, temporally coherent video sequences. Early works [singer2022make, Ho2022ImagenVH] extended 2D UNet-based image diffusion into 3D UNets for spatio-temporal generation, while latent diffusion approaches [he2211latent, rombach2022high, wang2023modelscope, chen2024videocrafter2] improved efficiency by operating in compressed latent spaces. More recent DiT-based architectures [liu2024sora, yang2024cogvideox, kong2024hunyuanvideo, wan2025wan] represent videos as sequences of latent tokens, enabling long-range temporal reasoning and disentangled modeling of appearance and motion. Building on these advances, we employ a pre-trained video diffusion model as a video-native prior within NVC. Instead of initializing diffusion from Gaussian noise as in video generation, GNVC-VD performs refinement directly on decoded spatio-temporal latents, learning a correction term that compensates for compression-induced degradation.
@@ 3 Proposed Method
In this work, we aim to achieve perceptually faithful and temporally coherent video reconstruction under extreme compression (<<0.03 bpp). Unlike prior generative codecs [ma2025diffusion, qi2025generative] that operate at the frame level with image-based priors [vqgan, stablediffusion], our GNVC-VD processes entire video sequences using a pre-trained video diffusion model [wan2025wan], enabling joint spatio-temporal latent compression and refinement across intra- and inter-frames. This design effectively captures long-range temporal dependencies and restores fine details while mitigating flickering artifacts. We first outline the overall framework (Section 3.1), followed by the spatio-temporal latent compression (Section 3.2), the diffusion-based latent refinement (Section 3.3), and the two-stage training strategy (Section 3.4).
@@ 3.1 Framework of GNVC-VD
As illustrated in Fig. 3(a), given an input video $V\in\mathbb{R}^{(1+T)\times H\times W\times 3}$, a 3D causal VAE encoder $\mathcal{E}$ from Wan2.1 [wan2025wan] encodes it into a compact spatio-temporal latent sequence:
@@
The reconstructed latent sequence $\boldsymbol{x}_{c}=\{\hat{l}_{t}\}_{t=1}^{1+T/4}$ is refined using the pre-trained VideoDiT from Wan2.1 [wan2025wan]. Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to obtain the noisy initialization $\boldsymbol{x}_{t_{N}}=\boldsymbol{x}_{c}+\epsilon$, which is iteratively denoised under the flow-matching formulation to generate the refined latent sequence $\tilde{\boldsymbol{x}}_{1}=\{\tilde{l}_{t}\}_{t=1}^{1+T/4}$:
@@
$\mathcal{E},\mathcal{D}$ 3D causal encoder / decoder (Wan2.1 [wan2025wan])
@@ 3.2 Spatio-Temporal Latent Compression
For the anchor latent $l_{1}$ (corresponding to the I-frame), we apply a separate transform coding module [ELIC] without temporal conditioning to initialize the sequence. For predictive latents $\{l_{t}\}_{t>1}$, each $l_{t}$ is encoded conditioned on the previously decoded latent $\hat{l}_{t-1}$ to reduce temporal redundancy, following the design philosophy of DCVC-RT [DCVC-RT]. A temporal context feature $f_{t-1}$ is extracted from $\hat{l}_{t-1}$ and injected into both the analysis transform $g_{a}$ and the synthesis transform $g_{s}$, extending Eq. (2) to:
@@ 3.3 Flow-Matching Latent Refinement
To further enhance perceptual quality at extremely low bitrates, GNVC-VD introduces a unified latent refinement module that leverages the pre-trained VideoDiT as a powerful video-native prior. Unlike prior diffusion-enhanced codec [qi2025generative] that refine frame-wise latents independently, our method performs refinement directly in the 3D latent space and jointly enhances the entire sequence of I- and P-frame latents, ensuring spatio-temporal coherence and texture consistency.
Preliminary on Flow Matching. Recent VideoDiT architectures adopt a flow-matching formulation [kong2024hunyuanvideo, wan2025wan] to train the diffusion models, which formulates generative modeling as learning a continuous velocity field $\boldsymbol{v}_{\tau}$ that transports a noisy sample $\boldsymbol{x}_{0}\!\sim\!\mathcal{N}(0,\boldsymbol{I})$ toward the data manifold $\boldsymbol{x}_{1}$. Given a probability path $\boldsymbol{x}_{\tau}$, the model predicts the instantaneous velocity $\boldsymbol{v}_{\tau}=\frac{d\boldsymbol{x}_{\tau}}{d\tau}$ that aligns with the target flow from $\boldsymbol{x}_{0}$ to $\boldsymbol{x}_{1}$, allowing deterministic generation and partial denoising without stochastic sampling.
@@
Figure 4: Rate–distortion curves on the HEVC-B [flynn16common], UVG [UVG], and MCL-JCV [MCL-JCV] in the ultra-low bitrate regime (<0.03<0.03 bpp). We report perceptual quality in terms of LPIPS and DISTS in the ultra-low bitrate regime (<0.03<0.03 bpp). GNVC-VD consistently achieves the best perceptual quality, clearly outperforming traditional codecs (HEVC, VVC), learned codecs (DCVC-FM, DCVC-RT), and generative baselines (GLC-Video).
@@ 4 Experiments
Figure 6: Visual comparison of temporal consistency. Ground-truth frames at t=0,1,5t{=}0,1,5 are shown on the left. On the right, GLC-Video [qi2025generative] displays clear temporal flickering—textures drift and vary across frames—while GNVC-VD produces stable, temporally coherent reconstructions.
@@ 4.1 Experimental Setup
Datasets. For training, we use the Vimeo-90k dataset [Vimeo90k] with 5-frame clips to pre-train GNVC-VD, and extend the original Vimeo videos [ori_vimeo] into longer 25-frame sequences for fine-tuning. Evaluation is conducted on widely used benchmarks, including HEVC Class B [flynn16common], UVG [UVG], and MCL-JCV [MCL-JCV].
Implementation Details. We adopt Wan2.1 [wan2025wan] as the pre-trained video diffusion model. The contextual transform coding network follows the architecture of DCVC-RT [DCVC-RT], and we use conditioning adapter blocks similar to those in VACE [vace]. In Stage I, GNVC-VD is trained on 5-frame Vimeo-90k clips using $256\times 256$ patches, a batch size of 8, and a learning rate of $1\times 10^{-4}$ for 40k iterations with $\lambda_{r}=0.5$. We then fine-tune on longer Vimeo sequences by resizing the shorter side to 512 pixels and randomly cropping $256\times 256$ patches. The model is trained for 30k iterations on 9-frame clips and another 30k on 13-frame clips, both with a learning rate of $1\times 10^{-4}$. In Stage II, we fine-tune for an additional 100k iterations with $\lambda_{r}\in\{0.05,0.1,0.25,0.5\}$ and a batch size of 2, while progressively reducing the learning rate from $5\times 10^{-5}$ to $1\times 10^{-5}$. The partial noise level is fixed at $t_{N}=0.7$, and the number of flow refinement steps $L$ is set to 5. All experiments are conducted on two NVIDIA A800 GPUs. Additional implementation details are provided in Appendix Section A.
@@
Compared Methods. We compare GNVC-VD against several state-of-the-art video compression approaches spanning traditional hybrids (HEVC [sullivan2012overview] and VVC [bross2021overview] ), neural codecs (DCVC-FM [DCVC-FM] and DCVC-RT [DCVC-RT]), and recent generative compression models (PLVC [plvc] and GLC-Video [qi2025generative]). All learned baselines are evaluated using official implementations or author-reported results for fair comparison. Following prior NVC studies [DCVC, DCVC-TCM, DCVC-HEM, DCVC-DC, DCVC-FM, DCVC-RT], we evaluate the first 96 frames of each test sequence under a low-delay prediction configuration. Since the video diffusion prior is pre-trained in RGB space, all baselines are tested under the same RGB mode to ensure a fair comparison protocol.
Evaluation Metrics. We evaluate GNVC-VD along three dimensions: perceptual quality, compression efficiency, and temporal consistency. Perceptual quality is measured using LPIPS-VGG [LPIPS] and DISTS [DISTS], where the VGG-based LPIPS variant is used as it aligns better with subjective perception in generative compression. Compression efficiency is quantified by Bits Per Pixel (BPP). Temporal consistency is assessed using CLIP-F [radford2021learning] and the warp error $E_{\text{warp}}$ [lai2018learning], where CLIP-F evaluates semantic continuity across frames, and $E_{\text{warp}}$ measures low-level alignment by comparing $\hat{X}_{t}$ with the flow-warped $\hat{X}_{t+1}$ as $E_{\text{warp}}=\big(\sum_{i}M_{t}^{i}\big)^{-1}\sum_{i}M_{t}^{i}\|\hat{X}_{t}^{i}-W(\hat{X}_{t+1},F_{t\rightarrow t+1})^{i}\|_{2}^{2}$, with $F_{t\rightarrow t+1}$ estimated by RAFT [teed2020raft] and $M_{t}$ denoting the non-occlusion mask. Additional metrics and results are provided in Appendix Section C.1.
Table 3: BD-Rate (%) comparisons anchoring by VVC [bross2021overview].

### v2

@@ Abstract
Figure 1: Qualitative comparison on ultra-low bitrate video compression. Traditional and learned codecs produce blurry frames. Generative approaches such as GLC-Video [38] yield sharper textures but introduce structural hallucinations and unstable details, causing pronounced temporal flickering (see Fig. 2). Leveraging a video-native diffusion prior, GNVC-VD produces coherent fine textures with strong temporal stability. Zoom in for best view.
@@ 1 Introduction
Neural video compression (NVC) [30, 15, 23, 43, 24, 25, 26, 16] has advanced rapidly in recent years, with learned codecs now surpassing traditional hybrid standards such as HEVC [45] and VVC [2] in rate–distortion optimization. However, when bitrate drops to the ultra-low regime, distortion-driven objectives (e.g., MSE) tend to oversmooth textures and erase fine structures, causing a sharp decline in perceptual realism. Improving perceptual quality under such extreme compression remains an open and fundamental challenge for NVC.
@@
In the image domain, this perceptual collapse has been largely alleviated. Recent generative image codecs [32, 50, 17, 3, 28, 36, 55, 19] leverage large pre-trained GANs [9, 7] or diffusion models [42] to recover high-frequency textures, producing visually convincing reconstructions even at extremely low bitrates. This naturally raises the question: can the same strategy be extended to video compression?
@@
Unfortunately, videos impose a much stricter requirement—temporal coherence. Although recent perceptual video codecs [31, 38] integrate image generative priors [7, 42], such priors remain inherently static and lack any modeling of temporal dynamics. As a result, codecs built upon them remain fundamentally frame-level: even with adjacent-frame conditioning, the generative prior cannot capture long-range temporal structure. Consequently, the restored appearance drifts over time, leading to the well-known perceptual flickering that becomes especially severe at ultra-low bitrates, as illustrated in Fig. 2.
@@
Recently, video diffusion models (VDMs)—especially those based on diffusion transformers (DiTs) [29, 53, 20, 47]—offer a natural path forward. Trained on large-scale video data, they learn spatio-temporal latent representations that capture appearance, motion, and long-range dependencies within a unified structure, enabling the synthesis of sequences with coherent texture and motion. These properties make VDMs an ideal generative prior for video compression, motivating us to rethink decoding not as independent frame reconstruction but as a sequence-level conditional denoising process guided by a video-native model.
@@
Building on this insight, we introduce GNVC-VD, the first generative NVC framework that fully leverages a pre-trained video diffusion transformer (VideoDiT). Unlike prior perceptual codecs [31] constrained by image generative priors and thus limited to frame-wise enhancement, GNVC-VD redesigns the entire coding pipeline around sequence-level compression and generative refinement, enabling the diffusion prior to guide reconstruction beyond frame-wise prediction. At its core, GNVC-VD integrates two tightly coupled components: (1) a conditional contextual transform codec that compresses the spatio-temporal latent representations while preserving long-range temporal structure, and (2) a flow-matching–based latent refinement module that performs sequence-level generative denoising across both intra- and inter-frame latents, driven by the video DiT.
@@
Figure 2: (a) Spatial and ttxx comparisons. Traditional and learned codecs lose fine textures, while GLC-Video [38] exhibits sharp but unstable structures that cause temporal flickering. GNVC-VD preserves clean textures and stable motion. (b) Frame-wise warp error $E_{\text{warp}}$ further confirms GNVC-VD’s temporal stability, in contrast to the large fluctuations of GLC-Video.
@@ 2 Related Works
Neural Video Compression [30, 15, 40, 10, 23, 43, 24, 25, 26, 16, 31, 38] has made substantial progress in recent years, achieving strong RD performance across a wide range of bitrates. By learning compact latent representations and temporal dependencies in an end-to-end manner, NVC methods outperform traditional hybrid codecs in both PSNR and MS-SSIM metrics. Prior works can be broadly categorized into residual-based [30, 15], 3D autoencoder-based [40, 10], and conditional coding-based architectures [23, 43, 24, 25, 26, 16]. Among these, conditional coding approaches—such as the DCVC family [23, 43, 24, 25, 26, 16]—have set new state-of-the-art results by using decoded features as context to guide motion estimation, latent prediction, and entropy modeling. However, as these methods are primarily optimized for distortion-oriented objectives (e.g., MSE), their reconstructions tend to be overly smooth and lack fine textures at extremely low bitrates. This reveals the inherent limitation of current NVC frameworks and calls for perceptually optimized compression to better preserve visual realism under extreme bitrate constraints.
@@
Perceptual Compression with Generative Prior. To enhance perceptual quality at ultra-low bitrates, recent studies introduce generative priors—termed generative compression [1, 34, 36, 32, 50, 17, 22, 3, 28, 27, 55, 19, 38, 31]—which leverage learned generative models to guide reconstruction and recover realistic textures beyond pixel fidelity. In the image domain, early works incorporate adversarial losses [1, 34, 36] or VQ-based tokenization [32, 50, 17] to achieve perceptually convincing reconstructions under extreme rate constraints. Diffusion-based frameworks [22, 3, 28, 27, 55, 19] further improve perceptual quality by reformulating decoding as conditional denoising guided by compact latents and leveraging large-scale text-to-image diffusion models prior [42].
@@
Extending this paradigm to videos, recent methods such as GLC-Video [38] and DiffVC [31] adapt pre-trained image generative models by either encoding frame-wise generative embeddings or applying diffusion-based frame enhancement. However, these approaches still depend on image-domain priors without explicit temporal modeling, leading to flickering and motion inconsistency at ultra-low bitrates, as illustrated in Fig. 2(b). In contrast, our work is the first to introduce a video generative prior into NVC, jointly encoding spatio-temporal latents and performing unified latent refinement to achieve temporally coherent and perceptually realistic reconstructions under extreme compression.
@@
Video Diffusion Models have emerged as powerful generative frameworks capable of synthesizing high-quality, temporally coherent video sequences. Early works [44, 14] extended 2D UNet-based image diffusion into 3D UNets for spatio-temporal generation, while latent diffusion approaches [13, 41, 49, 4] improved efficiency by operating in compressed latent spaces. More recent DiT-based architectures [29, 53, 20, 47] represent videos as sequences of latent tokens, enabling long-range temporal reasoning and disentangled modeling of appearance and motion. Building on these advances, we employ a pre-trained video diffusion model as a video-native prior within NVC. Instead of initializing diffusion from Gaussian noise as in video generation, GNVC-VD performs refinement directly on decoded spatio-temporal latents, learning a correction term that compensates for compression-induced degradation.
@@ 3 Proposed Method
In this work, we aim to achieve perceptually faithful and temporally coherent video reconstruction under extreme compression (<<0.03 bpp). Unlike prior generative codecs [31, 38] that operate at the frame level with image-based priors [7, 42], our GNVC-VD processes entire video sequences using a pre-trained video diffusion model [47], enabling joint spatio-temporal latent compression and refinement across intra- and inter-frames. This design effectively captures long-range temporal dependencies and restores fine details while mitigating flickering artifacts. We first outline the overall framework (Section 3.1), followed by the spatio-temporal latent compression (Section 3.2), the diffusion-based latent refinement (Section 3.3), and the two-stage training strategy (Section 3.4).
@@ 3.1 Framework of GNVC-VD
As illustrated in Fig. 3(a), given an input video $V\in\mathbb{R}^{(1+T)\times H\times W\times 3}$, a 3D causal VAE encoder $\mathcal{E}$ from Wan2.1 [47] encodes it into a compact spatio-temporal latent sequence:
@@
The reconstructed latent sequence $\boldsymbol{x}_{c}=\{\hat{l}_{t}\}_{t=1}^{1+T/4}$ is refined using the pre-trained VideoDiT from Wan2.1 [47]. Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to obtain the noisy initialization $\boldsymbol{x}_{t_{N}}=\boldsymbol{x}_{c}+\epsilon$, which is iteratively denoised under the flow-matching formulation to generate the refined latent sequence $\tilde{\boldsymbol{x}}_{1}=\{\tilde{l}_{t}\}_{t=1}^{1+T/4}$:
@@
$\mathcal{E},\mathcal{D}$ 3D causal encoder / decoder (Wan2.1 [47])
@@ 3.2 Spatio-Temporal Latent Compression
For the anchor latent $l_{1}$ (corresponding to the I-frame), we apply a separate transform coding module [11] without temporal conditioning to initialize the sequence. For predictive latents $\{l_{t}\}_{t>1}$, each $l_{t}$ is encoded conditioned on the previously decoded latent $\hat{l}_{t-1}$ to reduce temporal redundancy, following the design philosophy of DCVC-RT [16]. A temporal context feature $f_{t-1}$ is extracted from $\hat{l}_{t-1}$ and injected into both the analysis transform $g_{a}$ and the synthesis transform $g_{s}$, extending Eq. (2) to:
@@ 3.3 Flow-Matching Latent Refinement
To further enhance perceptual quality at extremely low bitrates, GNVC-VD introduces a unified latent refinement module that leverages the pre-trained VideoDiT as a powerful video-native prior. Unlike prior diffusion-enhanced codec [38] that refine frame-wise latents independently, our method performs refinement directly in the 3D latent space and jointly enhances the entire sequence of I- and P-frame latents, ensuring spatio-temporal coherence and texture consistency.
Preliminary on Flow Matching. Recent VideoDiT architectures adopt a flow-matching formulation [20, 47] to train the diffusion models, which formulates generative modeling as learning a continuous velocity field $\boldsymbol{v}_{\tau}$ that transports a noisy sample $\boldsymbol{x}_{0}\!\sim\!\mathcal{N}(0,\boldsymbol{I})$ toward the data manifold $\boldsymbol{x}_{1}$. Given a probability path $\boldsymbol{x}_{\tau}$, the model predicts the instantaneous velocity $\boldsymbol{v}_{\tau}=\frac{d\boldsymbol{x}_{\tau}}{d\tau}$ that aligns with the target flow from $\boldsymbol{x}_{0}$ to $\boldsymbol{x}_{1}$, allowing deterministic generation and partial denoising without stochastic sampling.
@@
Figure 4: Rate–distortion curves on the HEVC-B [8], UVG [35], and MCL-JCV [48] in the ultra-low bitrate regime (<0.03<0.03 bpp). We report perceptual quality in terms of LPIPS and DISTS in the ultra-low bitrate regime (<0.03<0.03 bpp). GNVC-VD consistently achieves the best perceptual quality, clearly outperforming traditional codecs (HEVC, VVC), learned codecs (DCVC-FM, DCVC-RT), and generative baselines (GLC-Video).
@@ 4 Experiments
Figure 6: Visual comparison of temporal consistency. Ground-truth frames at t=0,1,5t{=}0,1,5 are shown on the left. On the right, GLC-Video [38] displays clear temporal flickering—textures drift and vary across frames—while GNVC-VD produces stable, temporally coherent reconstructions.
@@ 4.1 Experimental Setup
Datasets. For training, we use the Vimeo-90k dataset [51] with 5-frame clips to pre-train GNVC-VD, and extend the original Vimeo videos [37] into longer 25-frame sequences for fine-tuning. Evaluation is conducted on widely used benchmarks, including HEVC Class B [8], UVG [35], and MCL-JCV [48].
Implementation Details. We adopt Wan2.1 [47] as the pre-trained video diffusion model. The contextual transform coding network follows the architecture of DCVC-RT [16], and we use conditioning adapter blocks similar to those in VACE [18]. In Stage I, GNVC-VD is trained on 5-frame Vimeo-90k clips using $256\times 256$ patches, a batch size of 8, and a learning rate of $1\times 10^{-4}$ for 40k iterations with $\lambda_{r}=0.5$. We then fine-tune on longer Vimeo sequences by resizing the shorter side to 512 pixels and randomly cropping $256\times 256$ patches. The model is trained for 30k iterations on 9-frame clips and another 30k on 13-frame clips, both with a learning rate of $1\times 10^{-4}$. In Stage II, we fine-tune for an additional 100k iterations with $\lambda_{r}\in\{0.05,0.1,0.25,0.5\}$ and a batch size of 2, while progressively reducing the learning rate from $5\times 10^{-5}$ to $1\times 10^{-5}$. The partial noise level is fixed at $t_{N}=0.7$, and the number of flow refinement steps $L$ is set to 5. All experiments are conducted on two NVIDIA A800 GPUs. Additional implementation details are provided in Appendix Section A.
@@
Compared Methods. We compare GNVC-VD against several state-of-the-art video compression approaches spanning traditional hybrids (HEVC [45] and VVC [2] ), neural codecs (DCVC-FM [26] and DCVC-RT [16]), and recent generative compression models (PLVC [52] and GLC-Video [38]). All learned baselines are evaluated using official implementations or author-reported results for fair comparison. Following prior NVC studies [23, 43, 24, 25, 26, 16], we evaluate the first 96 frames of each test sequence under a low-delay prediction configuration. Since the video diffusion prior is pre-trained in RGB space, all baselines are tested under the same RGB mode to ensure a fair comparison protocol.
Evaluation Metrics. We evaluate GNVC-VD along three dimensions: perceptual quality, compression efficiency, and temporal consistency. Perceptual quality is measured using LPIPS-VGG [54] and DISTS [6], where the VGG-based LPIPS variant is used as it aligns better with subjective perception in generative compression. Compression efficiency is quantified by Bits Per Pixel (BPP). Temporal consistency is assessed using CLIP-F [39] and the warp error $E_{\text{warp}}$ [21], where CLIP-F evaluates semantic continuity across frames, and $E_{\text{warp}}$ measures low-level alignment by comparing $\hat{X}_{t}$ with the flow-warped $\hat{X}_{t+1}$ as $E_{\text{warp}}=\big(\sum_{i}M_{t}^{i}\big)^{-1}\sum_{i}M_{t}^{i}\|\hat{X}_{t}^{i}-W(\hat{X}_{t+1},F_{t\rightarrow t+1})^{i}\|_{2}^{2}$, with $F_{t\rightarrow t+1}$ estimated by RAFT [46] and $M_{t}$ denoting the non-occlusion mask. Additional metrics and results are provided in Appendix Section C.1.
Table 3: BD-Rate (%) comparisons anchoring by VVC [2].